Cheryl wants to put a border around her window. The window is 3 feet wide and 5 feet high. How much border does she need to go around the window?
Answer by Alan3354(69443)

The perimeter of the window is 2W + 2L = 2*3 + 2*5 = 16 feet
-------------
"How much" is not clear. Do you mean the lenght, or the area?